You plan to do a expedition to Patagonia during Spring Break in 3 years. The trip will cost $9,000. You already have $3,000 in the bank and you plan to make a deposit every year beginning one year from today for the next 3 years. If your bank pays 6% EAR, how much does the yearly deposit need to be if you are to have $9,000 3 years from today?

A) $1,705 B) $1,712 C) $1,885 D) $2,245

The correct answer is A) $1,705 but I need to know the steps for solving this on a financial calculator
Answer #1

Amount required in 3 years = $9,000

Future value of amount already in Bank = Present value*(1+Interest Rate)^Number of years

= 3000(1+6%)^3

= $3,573.05

Additional amount required = 9000-3573.05 = $5,426.95

Let annual deposits be x

Future value of annuity = Annual amount*[{(1+r)^n -1 }/r]

5426.95 = x*[{(1.06)^3 - 1}/0.06]

5426.95 = 3.1836x

x = $1,704.66

Hence, the answer is A. $1,705